OpenStudy (anonymous):

A bag contains 60 marbles of various colors. Shana draws 48 times, each time replacing the marble before drawing again. She gets a total of 8 green marbles. Which is best prediction of how many green marbles are in the bag? A. 4 B. 6 C. 8 D. 10

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

can u anwser it

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

who ever gets it right gets a medal

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Every time you draw a marble, you have a 1 in 60 chance of picking any particular marble. If 8 of the 48 times you draw a marble are green, then statistically you draw 1 green marble per 6 draws. 8/48=1/6. Then if 1/6 of the total marbles are green, you can find the total number of green marbles by multiplying this proportion by the total number of marbles. green marbles = total marbles * % green marbles = 60 * 1 / 6 = 10